OpenStudy (anonymous):

Jim stood on flat ground 42 feet from the base of a tower. He measured the angle of elevation to the top of the tower as 73°40'. How tall is the tower?

OpenStudy (raden):

40' = (40/60)° = 0.667° 

OpenStudy (raden):

use defined of tan tan73.667° = x/42 solve for x

OpenStudy (raden):

with x is tall of the tower

OpenStudy (raden):

opppssss.. the sign of right angle is left :)